Hali’s current stock price is $36.00, its last dividend was $2.40, and its required rate of return is 12%. If dividends are expected to grow at a constant rate, g, in the future, and if the required rate of return is expected to remain at 12%, what is Hali’s expected stock price 5 years from now?
Answer #1

Required return=(D1/Current price)+Growth rate

0.12=(2.4*(1+Growth rate)/36+Growth rate

0.12=(2.4+2.4*Growth rate)/36+Growth rate

0.12=0.0667+0.0667*Growth rate+Growth rate

Growth rate=(0.12-0.0667)/(0.0667+1)

=5%

P5=Current price*(1+Growth rate)^5

=$36*(1.05)^5

=$45.95(Approx).